OpenStudy (sleepyhead314):

please help what is this magic? \(\frac{dy}{dt} = (y+2)(y-1)\)

OpenStudy (sleepyhead314):

@518nad

OpenStudy (sleepyhead314):

\(\frac{dy}{dt} = (y+2)(y-1)\)

OpenStudy (sleepyhead314):

\(\frac{dy}{dt} = y^2 + y -2\)

OpenStudy (518nad):

:) separate the weed from the good

OpenStudy (518nad):

no no it was fine the way it was before silly

OpenStudy (sleepyhead314):

yeah, I got stuck at u sub u = y + 1/2

OpenStudy (518nad):

okay so put all the ys on one side with dy and dt on oone side

OpenStudy (sleepyhead314):

\(\frac{1}{(y+2)(y-1)} dy= 1 dt \)

OpenStudy (518nad):